# Cryptocurrency as an Investable Asset Class: Coming of Age

### v2

@@ 3.7 Fact 7: In crypto, the (block)chain drives the gain
Empirically, Liu et al. (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices. Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns. In time-series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
@@
Liu et al. (2021) also examine how markets absorb this information and what conditions amplify its relevance. Event-time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre-announcement behavior and no post-announcement drift—in contrast to equities. The result gives evidence that investors fully incorporate the information at the weekly frequency.
@@
Lastly, Liu et al. (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in Fact 4.
3.8 Fact 8: Young cryptocurrency markets, old inefficiencies.
@@
Fact 8: The young cryptocurrency markets are still inefficient. But the origins are canonical: market segmentation and restrictions on capital flows. Cross-exchange cryptocurrency “arbitrage” is no free lunch.
@@
A central concept in finance is the law of one price—identical securities should sell for identical prices. The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978, Harrison & Kreps 1979). Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
@@
One of the main features of cryptocurrency is that it is possible to trade coins simultaneously on multiple exchanges and against a number of fiat and crypto pairs. For example, an investor could buy Bitcoin on Binance or Kraken, two prominent centralized exchanges, against fiat currencies such as the U.S. dollar or the euro, or even against another cryptocurrency such as Ethereum or Tether, a stablecoin. Makarov & Schoar (2020) and Borri & Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs. We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Bitcoin price discounts are computed each week relative to the dollar price on Kraken (the benchmark exchange). All quotes are converted to U.S. dollars using daily spot FX rates. In the 2014–2025 sample, the 10th and 90th percentiles of Bitcoin price discounts are –1.8% and 3.7%, respectively. These figures remain virtually unchanged in the post-2020 period (–2.1% and 3.5%). Moreover, discounts are highly persistent, with a half-life of roughly 1.1 weeks in both samples. These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage. [^13] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
@@
Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies. Each week, we sort BTC–fiat pairs by their discount and form five portfolios. The CARB return for a given pair is defined by buying Bitcoin on Kraken in week $t$ and selling it on the target pair in week t+1t{+}1; portfolio returns are the simple average across constituent pairs. Using weekly formation and holding periods substantially alleviates execution and timing concerns. In practice, capital controls, withdrawal limits, transaction costs, settlement latency, and venue-specific frictions prevent the cross-market trades needed to enforce the law of one price. Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly. Specifically, Borri & Shakhnov (2022) shows that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe. That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
@@
As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades. [^14] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman Fried. However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
@@
We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish łoty, Swiss franc, British pound, and U.S. dollar). These are the pairs that are realistically investable. Table 6, Panel B, documents that excess returns remain statistically significant but are orders of magnitude smaller: the long–short weekly return declines from about $0.68$ to roughly $0.038$ in the full sample, and measures around $0.041$ in the post-2020 sample. In other words, where the strategy is actually implementable, the apparent “free lunch” largely vanishes. These returns are not pure arbitrage—they are compensation for risk. Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu et al. (2023), is the most visible manifestation of the results described in this fact.
@@
3.9 Fact 9: When the funding dries up, we finally learn the worth of futures
Fact 9: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973, Merton 1973, Black 1976).
There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs. Among all derivatives, perpetual futures are the most common derivative contract in cryptocurrency markets. For example, Schmeling et al. (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample. Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot. Ackerer et al. (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures. The crypto-carry trade strategy of Schmeling et al. (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
@@ 3.9 Fact 9: When the funding dries up, we finally learn the worth of futures
The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis). The Carry Trade is defined as in Schmeling et al. (2023). The sample goes from August 1, 2020 to May 31, 2025.
3.10 Fact 10: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
Fact 10: As the market matures, there is growing disclosure and transparency in cryptocurrencies.
A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently. When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970). Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983, Dye 1985, Healy & Palepu 2001, Verrecchia 2001). However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001, Ma & Shleifer 2025). Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information. Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts & Zimmerman 1983).
@@ 3.10 Fact 10: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021, Liu et al. 2025, Du et al. 2025). Howell et al. (2020) and Bourveau et al. (2022) show the importance of voluntary disclosures in the initial coin offering stage. Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al. 2022, Bourveau et al. 2024).
@@ 4 ADDITIONAL DISCUSSIONS
The ten facts summarize key developments in cryptocurrencies through the lens of empirical asset pricing. In this section, we provide additional discussions on emerging topics, including retail investor trading in the cryptocurrency market, the non-fungible token market, and studies on the dark side of cryptocurrency usage.
@@
Recent research increasingly examines individual investor beliefs and behavior in cryptocurrency markets and compares and contrast them with those in traditional investments. Kogan et al. (2024) show that the trading behaviors of retail investors are drastically different between cryptocurrencies and traditional assets, such as stocks and gold. In particular, retail investors are contrarian in stocks and gold, but trend-chasing in cryptocurrencies. Kogan et al. (2024) conjecture that the trend-chasing behavior is due to cryptocurrency price changes affecting the likelihood of future widespread adoption, which leads retail investors to update their price expectations. Chan et al. (2020) use data from a medium-sized cryptocurrency exchange and document that investor portfolios are typically small in value, highly concentrated in a few coins, and characterized by short holding periods. Weber et al. (2023) complement this evidence using repeated large-scale surveys of U.S. households, showing that cryptocurrency investors differ systematically from other investors. Alvarez et al. (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption. Han et al. (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets. Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies. They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
@@
The bulk of this review is about fungible tokens. A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs). Goetzmann & Nozari (2022) is the first to construct repeated sales return indices of NFTs and examine their properties. Borri et al. (2022a) provide a comprehensive analysis of the NFT market. They highlight the importance of visual features in cryptocurrency valuation and how it leads to a conspicuous consumption effect in the market. Furthermore, Borri et al. (2022a) shows that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market. Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing. Oh et al. (2025) document preference-driven herding in NFT primary market sales.
@@
A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years. Foley et al. (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley et al. (2019) are largely driven by double-counting. Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation. Amiram et al. (2022) show that abnormal spikes in on-chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, and Cong et al. (2025) provide an anatomy of crypto-enabled crimes. Li et al. (2025) document pervasive cryptocurrency pump-and-dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre-event run-ups consistent with insider gains at outsiders’ expense.

### v3

@@ 3.4 Stylized fact 4: In crypto, the (block)chain drives the gain
Empirically, Liu, Tsyvinski & Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices. Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns. In time‑series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
@@
Liu, Tsyvinski & Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance. Event‑time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre‑announcement behavior and no post‑announcement drift—in contrast to equities. The result gives evidence that investors fully incorporate the information at the weekly frequency.
@@
Lastly, Liu, Tsyvinski & Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
Taken together, these results connect to classic theories of networks and information diffusion and suggest that network-based adoption and propagation forces are quantitatively important for cryptocurrency returns—arguably more so than in mature equity markets.
3.5 Stylized fact 5: Young cryptocurrency markets, old inefficiencies.
@@
Stylized fact 5: The young cryptocurrency markets are still inefficient. But the origins are canonical: market segmentation and restrictions on capital flows. Cross-exchange cryptocurrency “arbitrage” is no free lunch.
@@
A central concept in finance is the law of one price—identical securities should sell for identical prices. The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison & Kreps 1979). Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
One of the main features of cryptocurrency is that it is possible to trade coins simultaneously on multiple exchanges and against a number of fiat and crypto pairs. For example, an investor could buy Bitcoin on Binance or Kraken, two prominent centralized exchanges, against fiat currencies such as the U.S. dollar or the euro, or even against another cryptocurrency such as Ethereum or Tether, a stablecoin.
@@
Makarov & Schoar (2020) and Borri & Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs. We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Bitcoin price discounts are computed each week relative to the dollar price on Kraken (the benchmark exchange). All quotes are converted to U.S. dollars using daily spot FX rates. In the 2014–2025 sample, the 10th and 90th percentiles of Bitcoin price discounts are –1.8% and 3.7%, respectively. These figures remain virtually unchanged in the post-2020 period (–2.1% and 3.5%). Moreover, discounts are highly persistent, with a half-life of roughly 1.1 weeks in both samples. These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage.
Fragmentation and limits to arbitrage also have implications for market fragility and flash-crash dynamics. Menkveld & Yueshen (2019) study the May 6, 2010 U.S. equity Flash Crash and argue that a disruption in cross-market arbitrage can make fragmented markets fragile and result in price crashes; they also refer to analogous flash-crash episodes in cryptocurrency markets, including in Ethereum (June 21, 2017) and Bitcoin (October 10, 2017). Because our baseline analysis is conducted at the weekly frequency using close-to-close returns, these very short-lived intraday dislocations are less likely to affect our main results, though they remain relevant for studies of market microstructure and tail risk. [^17] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
@@
Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies. Each week, we sort BTC–fiat pairs by their discount and form five portfolios. The CARB return for a given pair is defined by buying Bitcoin on Kraken in week $t$ and selling it on the target pair in week t+1t{+}1; portfolio returns are the simple average across constituent pairs. Using weekly formation and holding periods substantially alleviates execution and timing concerns. In practice, capital controls, withdrawal limits, transaction costs, settlement latency, and venue-specific frictions prevent the cross-market trades needed to enforce the law of one price. Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly. Specifically, Borri & Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe. That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
@@
As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades. [^18] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman-Fried. However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
@@
We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S. dollar). These are the pairs that are realistically investable. Table 6, Panel B, documents that excess returns remain statistically significant but are orders of magnitude smaller: the long–short weekly return declines from about $0.68$ to roughly $0.038$ in the full sample, and measures around $0.041$ in the post-2020 sample. In other words, where the strategy is actually implementable, the apparent “free lunch” largely vanishes. These returns are not pure arbitrage—they are compensation for risk. Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov & Schoar (2023), is the most visible manifestation of the results described in this fact. More broadly, the evidence from cryptocurrency markets provides a clean illustration of the classic limits-to-arbitrage view: when trading, settlement, and capital-mobility frictions are economically large, violations of the law of one price can persist rather than being immediately arbitraged away.
@@
3.6 Stylized fact 6: When the funding dries up, we finally learn the worth of futures
Stylized fact 6: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973; Merton 1973; Black 1976).
There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs. Among all derivatives, perpetual futures are the most common derivative contracts in cryptocurrency markets. For example, Schmeling, Schrimpf & Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample. Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot. Ackerer, Hugonnier & Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures. The crypto-carry trade strategy of Schmeling, Schrimpf & Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
@@ 3.6 Stylized fact 6: When the funding dries up, we finally learn the worth of futures
The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis). The Carry Trade is defined as in Schmeling, Schrimpf & Todorov (2023). The sample goes from August 1, 2020 to May 31, 2025.
3.7 Stylized fact 7: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
Stylized fact 7: As the market matures, there is growing disclosure and transparency in cryptocurrencies.
A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently. When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970). Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy & Palepu 2001; Verrecchia 2001). However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma & Shleifer 2025). Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information. Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts & Zimmerman 1983).
@@
On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021; Liu, Sheng & Wang 2025; Du, Huang & Wu 2025). Howell, Niessner & Yermack (2020) and Bourveau et al. (2022) show the importance of voluntary disclosures in the initial coin offering stage. Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al. 2022; Bourveau, Brendel & Schoenfeld 2024).
@@ 4 ADDITIONAL DISCUSSIONS
The seven stylized facts summarize key developments in cryptocurrencies through the lens of empirical asset pricing. In this section, we provide additional discussions on emerging topics, including retail investor trading in the cryptocurrency market, the non-fungible token market, and studies on the dark side of cryptocurrency usage.
@@
Recent research increasingly examines individual investor beliefs and behavior in cryptocurrency markets and compares and contrast them with those in traditional investments. Kogan et al. (2024) show that the trading behaviors of retail investors are drastically different between cryptocurrencies and traditional assets, such as stocks and gold. In particular, retail investors are contrarian in stocks and gold, but trend-chasing in cryptocurrencies. Kogan et al. (2024) conjecture that the trend-chasing behavior is due to cryptocurrency price changes affecting the likelihood of future widespread adoption, which leads retail investors to update their price expectations. Chan et al. (2020) use data from a medium-sized cryptocurrency exchange and document that investor portfolios are typically small in value, highly concentrated in a few coins, and characterized by short holding periods. Weber et al. (2023) complement this evidence using repeated large-scale surveys of U.S. households, showing that cryptocurrency investors differ systematically from other investors. Alvarez, Argente & Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption. Han, Liu & Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets. Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies. They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
@@
The bulk of this review is about fungible tokens. A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs). Goetzmann & Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties. Borri et al. (2022a) provide a comprehensive analysis of the NFT market. They highlight the importance of visual features in cryptocurrency valuation and how it leads to a conspicuous consumption effect in the market. Furthermore, Borri et al. (2022a) show that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market. Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing. Oh, Rosen & Zhang (2025) document preference-driven herding in NFT primary market sales.
@@
A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years. Foley, Karlsen & Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley, Karlsen & Putniņš (2019) are largely driven by double-counting. Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation. Amiram, Jørgensen & Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al. (2025) provide an anatomy of crypto-enabled crimes and identify ransomware as the dominant activity, while Cong et al. (2023a) review crypto-related scams, including Ponzi schemes and “rug pulls” in decentralized projects and token offerings. Li, Shin & Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.